OpenStudy (anonymous):

a side of rhombus ABCD measures 100 feet. the measures of

OpenStudy (campbell_st):

start by drawing the diagram  so you know 2 sides and the included angle looks like the law of cosines is needed to find the length of AC \[b^2 = a^2 + c^2 - 2ac \times \cos(B)\] hope this helps

OpenStudy (anonymous):

thank you! you're the best! :)

OpenStudy (campbell_st):

oops the angle should be \[110^o 20'\] missed the 20 minutes...